OpenStudy (anonymous):

what is the equation of the quadratic graph with a focus of (3,6) and a directrix of y=4???

OpenStudy (anonymous):

@phi I don't understand what I am supposed to do. am I supposed to plug (3,6) into the vertex form y=a(x-h)^2+k

OpenStudy (phi):

I guess they want you to learn that the vertex is the point exactly in between the focus and the directrix. I always plot the focus and the directrix to make it more clear what is going on.

OpenStudy (phi):

can you find the vertex ?

OpenStudy (phi):

the other thing you need to know is "a" in y=a(x-h)^2+k is equal to \( \frac{1}{4p} \) "p" is the distance from the vertex to the focus (or the distance from the vertex to the directrix or ½ the distance between the focus and the directrix)

OpenStudy (phi):

Here is a plot

OpenStudy (anonymous):

so should i plug the numbers into the distance formula

OpenStudy (phi):

you don't need to. the shortest distance from a point to a line is a line that makes a right angle with the given line. in other words, the distance is the distance from the point *straight down* to the directrix

OpenStudy (anonymous):

so its just this??

OpenStudy (phi):

point B is not the closest point on the line y=4 y=4 is a line that goes horizontal. go *down* (not sideways or diagonally)

OpenStudy (phi):

yes. If you also plot the directrix y=4 it would look nice. the VERTEX will be on that line, exactly ½ between the 2 points.

OpenStudy (anonymous):

but the question says " what us the equation of the quadratic graph with a focus of (3,6) and directrix of y=4? I have the graph but i don't understand how to put it into an equation.

OpenStudy (phi):

The first step to answering the question is to find the vertex. The vertex is exactly half way between points A and B in your plot. can you find the x,y coordinates of the vertex?

OpenStudy (anonymous):

(3,5)

OpenStudy (phi):

yes, the vertex is at (3,5) that means you know h and k for your formula y = a(x-h)^2 + k

OpenStudy (phi):

remember (h,k) is the vertex. h is 3 and k is 5 for this problem

OpenStudy (anonymous):

what is a

OpenStudy (phi):

the other thing you need to know is "a" in y=a(x-h)^2+k "a" is equal to \( \frac{1}{4p} \) "p" is the distance from the vertex to the focus (or the distance from the vertex to the directrix or ½ the distance between the focus and the directrix)

OpenStudy (anonymous):

so a is 2

OpenStudy (anonymous):

or 1

OpenStudy (anonymous):

Someone help me!!

OpenStudy (anonymous):

Okay

OpenStudy (phi):

First what is "p" ?

OpenStudy (anonymous):

Which of the following does not describe the dependent variable? A. It is the range. B. It is the y-values. C. It is the output. D. It is the x-values. \

OpenStudy (phi):

"p" is the distance from the vertex to the focus the vertex is (3,5) and the focus is (3,6)

OpenStudy (anonymous):

so you ain't gonna help me?

OpenStudy (anonymous):

in a minute. first go make your own question. I am in the middle of getting help my self.

OpenStudy (phi):

you do not need to post a picture. what is the distance between point C (the vertex) and point A (the focus) ?

OpenStudy (anonymous):

Which values represent the independent variable? (–2, 4), (3, –2), (1, 0), (5, 5) A. {–2, 3, 1, 5} B. {4, –2, 0, 5} C. {–2, 4, 3, –2} D. {–2, –1, 0, 5}

OpenStudy (anonymous):

1

OpenStudy (phi):

you could use the distance formula \[ \sqrt{(3-3)^2 + (6-5)^2 } = \sqrt{0+1}= 1 \]

OpenStudy (phi):

now use a= 1/(4*p)

OpenStudy (phi):

Here is what your parabola looks like

OpenStudy (anonymous):

so its .25 or 1/4

OpenStudy (phi):

yes. now you have the full equation

OpenStudy (phi):

if you type that equation into geogebra, it will plot it.

OpenStudy (anonymous):

y=1/4(x-3)^2+5

OpenStudy (phi):

yes

OpenStudy (anonymous):

do you know how to change the geogebra graph to have all the lines on the coordinate plane

OpenStudy (phi):

The right most button up top. if you select it, it will move the center of the graph. If you click on its lower right corner (where the tiny triangle is) you can select zoom out.